OpenStudy (anonymous):

write the equation of a line with the given information: (4,8) and (16,32)

OpenStudy (shadowfiend):

(See above.)

OpenStudy (anonymous):

Find the slope which is 2. Use the point slope form to find the equation by using a point. y-y1 = m(x-x1) y - 8 = 2(x - 4) y = 2x - 8 + 8 y = 2x is the equation of the line